Problem 3. Use the Chinese Remainder Theorem to find all congruence classes that satisfy x2 = 1 mod 77.
Answer #1

5012 x² = 1 (mod 77) Then - x² = (mod 7) and x² = 1 (mod 11) Now, x² = 1 (mod 7) has only solt x = /(mod 7) similarly, x²= 1& 6 (modt) x= 1 (mod 11) (6 25 x= (6x2x11 + 1x8*7) (mod 77) 188 (mod 77) → 2 = x= 34 mod 77 (c) 2 = 1 (mod) & x= 10 (mod 11)